The length of a rectangle is 2in. more than twice its width. If the perimeter of the rectangle is 34 in., find the dimensions of the rectangle.
Thanks.

Answer by tutorcecilia(2152)

Perimeter = 2(length + width) [formula for the perimeter of a rectangle]
34=2(2w+w) [plug-in the values and simplify]
34/2=[2(2w+w)/2
17=2w+w
17=3w
17/3=(3w)/3
17/3=w
Therefore, the width = 17/3 in.
Length = 2w=2(17/3) = 34/3 = 11.333 in.
.
.
Check by plugging all of the values back into the original formula:
P=2(l+w)
34=2(2w+w)
34=2([(2)(17/3)+(17/3)]
34=2[34/3 + 17/3]
34=(68/3+34/3)
34=(102/3)
34=34 [checks out]